OpenStudy (anonymous):

How to take the derivative of f(x)= (x+2)^3, x=-1

OpenStudy (anonymous):

you want to take the derivative and then plug in the values x=-1 ?

OpenStudy (anonymous):

Yes but how

OpenStudy (anonymous):

power rule for this one \[f(x)=(x+2)^3\] \[f'(x)=3(x+2)^2\] finito

OpenStudy (anonymous):

Okay thanks i got it